### v1

# Merging LoRAs like Playing LEGO: Pushing the Modularity of LoRA to Extremes Through Rank-Wise Clustering

Abstract
Low-Rank Adaptation (LoRA) has emerged as a popular technique for fine-tuning large language models (LLMs) to various domains due to its modular design and widespread availability on platforms like Huggingface. This modularity has sparked interest in combining multiple LoRAs to enhance LLM capabilities. However, existing methods for LoRA composition primarily focus on task-specific adaptations that require additional training, and current model merging techniques often fail to fully leverage LoRA’s modular nature, leading to parameter interference and performance degradation. In this paper, we investigate the feasibility of disassembling and reassembling multiple LoRAs at a finer granularity, analogous to assembling LEGO blocks. We introduce the concept of Minimal Semantic Units (MSUs), where the parameters corresponding to each rank in LoRA function as independent units. These MSUs demonstrate permutation invariance and concatenation-summation equivalence properties, enabling flexible combinations to create new LoRAs. Building on these insights, we propose the LoRA-LEGO framework. This framework conducts rank-wise parameter clustering by grouping MSUs from different LoRAs into $k$ clusters. The centroid of each cluster serves as a representative MSU, enabling the assembly of a merged LoRA with an adjusted rank of $k$. Additionally, we apply a dual reweighting strategy to optimize the scale of the merged LoRA. Experiments across various benchmarks demonstrate that our method outperforms existing approaches in LoRA merging.
1 Introduction
Large Language Models (LLMs) like ChatGPT Achiam et al. (2023) and LLaMA Touvron et al. (2023) trained on vast amounts of general data, demonstrate remarkable performance in general tasks. To explore their potential for specialized tasks, adapting LLMs to specific domains by fine-tuning model parameters has become a critical area of research. In this context, Low-rank Adaptation (LoRA) Hu et al. (2021), as a parameter-efficient fine-tuning approach, has gained widespread recognition, also attributed to its modular design Liu et al. (2023); Yang et al. (2023b); Hadi et al. (2023). The modular nature of LoRA enables it to serve as plug-and-play plugins for LLMs, facilitating the storage and deployment of large collections of LoRAs on platforms like Hugging Face. The extensive availability of LoRAs has sparked considerable interest in combining multiple LoRAs into a unified adapter to significantly extend the capabilities of LLMs Yadav et al. (2024a); Xiao et al. (2024); Zhao et al. (2024b); Huang et al. (2023).
Previous methods for composing multiple LoRAs have primarily focused on assembling separate LoRAs tailored to specific downstream tasks, which generally require additional training Wu et al. (2023); Wang et al. (2024); Chronopoulou et al. (2023); Yadav et al. (2024a); Huang et al. (2023). Model merging Tang et al. (2024); Yadav et al. (2024b); Ilharco et al. (2022); Yang et al. (2024) offers an alternative approach by aggregating the parameters of multiple LoRAs into a unified adapter without extra training, producing a unified LoRA with more comprehensive capabilities. However, these methods typically employ element-wise parameter fusion, which can neglect and disrupt the internal semantic structure within LoRA. This disruption potentially leads to parameter interference (as discussed in §2.3), thereby hindering the performance of merged LoRA. This paper approaches LoRA merging from a novel perspective, focusing on the fine-grained modularization of LoRA by decomposing it into independent units, which enables the flexible reconstruction of a unified LoRA with comprehensive capabilities.
Figure 1: Further Modularization of LoRA: a) Each LoRA can be further modularized into multiple Minimal Semantic Units (MSUs), each corresponding to a column in ${\bm{A}}$ matrix and a row in matrix ${\bm{B}}$, differentiated by distinct colors. b) The MSUs within a LoRA display permutation invariance, implying that any rearrangement of the MSUs does not affect the output generated by the LoRA. c) Multiple LoRAs exhibit Concatenation-Summation Equivalence, indicating that the summation of outputs from various LoRAs is equivalent to the output of a singular LoRA constructed by concatenating their MSUs.
As illustrated in Fig.1, our motivation for further modularizing LoRA stems from the following insights: a) A rank of LoRA corresponds to a column in the down-projection matrix ${\bm{A}}$ and a row in the up-projection matrix ${\bm{B}}$. Therefore, we consider the parameters associated with each rank as a cohesive entity and define them as Minimal Semantic Units (MSUs), which serve as the fundamental building blocks of each LoRA. b) Within each LoRA, the MSUs exhibit the property of Permutation Invariance, indicating that any permutation of MSUs within a LoRA does not affect the adapter’s output. c) LoRA exhibits the Concatenation-Summation Equivalence property, which states that summing the outputs from multiple LoRAs is equivalent to the output of a single higher-ranked LoRA constructed by concatenating the MSUs of these LoRAs.
In this paper, we introduce a novel method called LoRA-LEGO, which is based on the insight that MSUs act as building blocks that form a LoRA and can be disassembled and reassembled like playing with LEGO. LoRA-LEGO consists of three main steps: (1) Grouping MSUs from candidate LoRAs into a MSU pool; (2) Clustering the MSU pool into $k$ clusters, where $k$ is the target rank of the merged LoRA; (3) Constructing the merged LoRA from the centroids of these clusters, with each centroid representing an MSU, thereby setting the merged LoRA’s rank to $k$. LoRA-LEGO enables the flexible combination of LoRAs with arbitrary ranks by clustering similar MSUs, at the same time effectively resolving parameter interference while merging. This approach allows for targeted rank adjustments in the merged LoRA to preserve task-specific knowledge.
We empirically validate the effectiveness of the proposed LoRA-LEGO in both Multi-task Tang et al. (2024) and Mixed-task Zhao et al. (2024b) scenarios. Experimental results show that LoRA-LEGO consistently outperforms other methods for LoRA merging, demonstrating notable flexibility and efficiency. Additionally, LoRA-LEGO can merge heterogeneous LoRAs of varying ranks, surpassing the capabilities of previous model merging methods. Moreover, it can also be applied to individual LoRAs for parameter pruning, revealing that retaining just 50% of the parameters can achieve performance comparable to the original model. Our contribution can be summarized as:
  • We investigate the modularization of LoRA, identifying the MSU as its fundamental building block, which is characterized by permutation invariance and concatenation-summation equivalence properties.
  • We introduce LoRA-LEGO that merges multiple LoRAs in a LEGO-like fashion by grouping, clustering, and reconstructing MSUs to seamlessly combine separate LoRAs.
  • Experimental results show that LoRA-LEGO can flexibly disassemble and reassemble LoRAs of any rank, surpassing other model merging methods in performance. Additionally, LoRA-LEGO can be effectively applied to individual LoRAs, enabling parameter pruning and a substantial reduction in LoRA parameters while maintaining comparable performance.
2 Preliminaries
2.1 Low-Rank Adaptation
Directly fine-tuning LLMs with full parameters is computationally intensive and is not feasible in low-resource scenarios. Based on the idea that only a small number of low-rank parameters need to be fine-tuned for sufficient performance in new domains, Hu et al. (2021) proposed the Low-Rank Adaptation, where the LoRA module can be combined with the pre-trained parameters in parallel for efficient inference.
Specifically, given pre-trained weights ${\bm{W}}_{0}\in\mathbb{R}^{d\times k}$ of a sub-module of LLM, the LoRA adds an extra trainable weight matrix as ${\bm{W}}_{0}+\Delta{\bm{W}}={\bm{W}}_{0}+{\bm{B}}{\bm{A}}$, where $\Delta{\bm{W}}$ can be decomposed into two smaller matrices ${\bm{B}}\in\mathbb{R}^{d\times r}$ and ${\bm{A}}\in\mathbb{R}^{r\times k}$, where $r$ stands for the rank of $\Delta{\bm{W}}$ and the rank $r\ll min(d,k)$. The forward pass for a layer ${\bm{y}}={\bm{W}}_{0}{\bm{x}}$ can be modified as follows:
${\bm{y}}={\bm{W}}_{0}{\bm{x}}+\Delta{\bm{W}}{\bm{x}}={\bm{W}}_{0}{\bm{x}}+{\bm{B}}{\bm{A}}{\bm{x}},$ (1)
where ${\bm{x}}\in\mathbb{R}^{d}$ is the input and the ${\bm{y}}\in\mathbb{R}^{d}$ denote the output.
2.2 Further modularization of LoRA
Before delving into the issue of LoRA merging, it is imperative to present several pivotal insights and definitions that could serve as fundamental components for constructing a LoRA module.
Definition 1.
Minimum Semantic Unit of LoRA. Let ${\bm{A}}$ and ${\bm{B}}$ be matrices in a LoRA module. For each index $i$, define the minimum semantic unit of LoRA as the combined vector ${\bm{s}}_{i}=[{\bm{a}}_{i},{\bm{b}}_{i}]$, where ${\bm{a}}_{i}$ is the $i$-th column of ${\bm{A}}$ and ${\bm{b}}_{i}$ is the $i$-th column of ${\bm{B}}^{T}$ (i.e., the transpose of the $i$-th row of ${\bm{B}}$).
In this context, each column of the down-projection matrix ${\bm{A}}$ and its corresponding row in the up-projection matrix ${\bm{B}}$ are treated as a cohesive unit, defined as a Minimum Semantic Unit (MSU). Each MSU contributes to a rank of the LoRA, encapsulating a distinct semantic fragment of the LoRA’s capacity. Through this definition, LoRAs exhibit the following properties.
Property 2.1.
Permutation Invariance. For a LoRA module parameterized by matrices ${\bm{A}}$ and ${\bm{B}}$, if the rows of ${\bm{A}}$ are permuted, then by performing a corresponding permutation of the columns of ${\bm{B}}$, the product of these matrices remains unchanged. Formally, let ${\bm{P}}$ be a permutation matrix that satisfies ${\bm{P}}^{T}{\bm{P}}={\bm{I}}$, where ${\bm{I}}$ is the identity matrix. If we permute the rows of ${\bm{A}}$ to obtain a new matrix ${\bm{A}}^{\prime}={\bm{P}}{\bm{A}}$, and correspondingly permute the columns of ${\bm{B}}$ to get ${\bm{B}}^{\prime}={\bm{B}}{\bm{P}}^{T}$, then, ${\bm{B}}{\bm{A}}={\bm{B}}^{\prime}{\bm{A}}^{\prime}$.
The property of permutation invariance indicates that the arrangement of MSUs within LoRA calculations can be altered without affecting LoRA’s output.
Property 2.2.
Concatenation-Summation Equivalence. Consider two LoRAs, $({\bm{A}}_{1},{\bm{B}}_{1})$ and $({\bm{A}}_{2},{\bm{B}}_{2})$, each of rank $r$. Specifically, matrices ${\bm{A}}_{1}$ and ${\bm{A}}_{2}$ are of size $\mathbb{R}^{r\times d}$, and ${\bm{B}}_{1}$ and ${\bm{B}}_{2}$ are of size $\mathbb{R}^{d\times r}$. Define the concatenated matrices as:
${\bm{A}}^{\prime}=\begin{bmatrix}{\bm{A}}_{1}\\ {\bm{A}}_{2}\end{bmatrix}\in\mathbb{R}^{2r\times d},\quad{\bm{B}}^{\prime}=\begin{bmatrix}{\bm{B}}_{1}&{\bm{B}}_{2}\end{bmatrix}\in\mathbb{R}^{d\times 2r}.$
The output vector ${\bm{y}}$ from the concatenated model is equivalent to the sum of the outputs from each individual LoRA model:
${\bm{y}}={\bm{B}}^{\prime}{\bm{A}}^{\prime}{\bm{x}}=({\bm{B}}_{1}{\bm{A}}_{1}+{\bm{B}}_{2}{\bm{A}}_{2}){\bm{x}}.$
Based on this property, we can synthesize the knowledge from all LoRAs by constructing a new LoRA through the concatenation of all MSUs from each LoRA. The computational result is equivalent to ensembling the outputs of all LoRAs. Based on these insights, we can draw the following conclusions:
Each LoRA can be modularized into multiple MSUs, with each MSU corresponding to a rank within the LoRA. These MSUs can be flexibly permuted and combined to construct a unified LoRA.
Figure 2: Two sources of parameter interference in LoRA merging. The left part illustrates how parameter misalignment can lead to interference; the right part demonstrates that knowledge conflict in merged LoRA layers can also result in parameter interference.
2.3 Problem Formulation and Challenges
Table 1: Performance degradation after merging misaligned LoRAs. “Original” refers to the performance of the unaltered LoRA, while “Misaligned” indicates the performance after merging the LoRA with a randomly permuted version of itself.
Task Original Misaligned
CoLA 61.63 60.96 (1.1% $\downarrow$)
MNLI 77.46 69.49 (10.3% $\downarrow$)
MRPC 68.00 68.50 (-0.7% $\downarrow$)
QNLI 77.25 60.44 (21.8% $\downarrow$)
QQP 75.83 66.94 (11.7% $\downarrow$)
RTE 52.22 54.44 (-4.2% $\downarrow$)
SST2 75.74 75.52 (0.3% $\downarrow$)
Overall 69.73 65.18 (5.74% $\downarrow$)
Consider a LLM denoted as $\mathcal{L}$ and a set of $p$ task-specific LoRAs, represented by $\Phi=\{\phi_{1},\phi_{2},\ldots,\phi_{p}\}$. Each LoRA $\phi_{i}$ is specialized for a particular task $\mathcal{T}_{i}$ and is crafted by incorporating low-rank matrices into different layers of $\mathcal{L}$, thereby tuning the model to better suit $\mathcal{T}_{i}$. For simplicity of notation, we denote the parameters of these low-rank matrices at any given layer for each LoRA $\phi_{i}$ as ${\bm{A}}_{i}$ and ${\bm{B}}_{i}$. The goal of merging these LoRAs is to synthesize a comprehensive LoRA $\phi^{\prime}$ that not only excels in all tasks encompassed by $\Phi$ but also generalizes well to unseen tasks. We discuss the difference between the LoRA merging setting and the previous model merging setting in the Appendix A.
A natural approach to performing LoRA merging involves a simple element-wise averaging of the parameters from each LoRA: $\phi^{\prime}=\frac{1}{p}\sum_{i=1}^{p}\phi_{i}$. However, parameter interference poses a significant challenge to effective LoRA merging. We identify two potential sources of parameter interference during LoRA merging and demonstrate through experiments that such interference can lead to performance degradation in the merged LoRA.
Table 2: Parameter interference due to knowledge conflict. “Tuning MSU” indicates the performance after tuning the added MSU for each task. “Avg MSU” denotes the performance achieved by directly merging these task-specific MSUs. “Concat MSU” represents the performance after concatenating these task-specific MSUs.
Task Tuning MSU Avg MSU Concat MSU
MNLI 86.17 46.24 (46.35$\downarrow$) 81.36 (5.58$\downarrow$)
MRPC 87.25 64.75 (25.78$\downarrow$) 81.25 (6.88$\downarrow$)
Overall 86.71 55.49 (36.06$\downarrow$) 81.31 (6.23$\downarrow$)
The first cause of parameter interference stems from parameter misalignment in LoRAs, as depicted in the left part of Fig.2. Accoding to Property 2.1, the MSUs of each LoRA can be permuted arbitrarily without affecting the functionality of the LoRA module. However, misalignment of MSU parameters when merging LoRAs can result in parameter interference. To investigate the impact of parameter misalignment on model performance, we conducted a controlled experiment using the Llama-2-7b model, training LoRAs on different tasks. For the parameters ${\bm{A}}$ and ${\bm{B}}$ of a task, we randomly generated a permutation matrix ${\bm{P}}$ and adjusted the parameters to ${\bm{A}}^{\prime}=({\bm{A}}+{\bm{P}}{\bm{A}})/2$ and ${\bm{B}}^{\prime}=({\bm{B}}+{\bm{B}}{\bm{P}}^{T})/2$. This adjustment simulates the merging of two identical LoRAs with misaligned parameters. The results, presented in Tab.1, indicate that parameter misalignment can lead to a decline in model performance, with some tasks experiencing significant performance degradation. Therefore, ideal merging entails alignming MSUs during LoRA merging to mitigate parameter interference.
Another source of parameter interference stems from knowledge conflict during LoRA merging. As depicted on the right side of Fig.2, knowledge conflict occurs when the merged LoRA lacks sufficient parameter space to encapsulate the comprehensive knowledge. This deficiency forces the merging of task-specific MSUs, resulting in parameter interference. To investigate the impact of knowledge conflict during LoRA merging, we conducted an experiment to demonstrate the performance degradation resulting from merging task-specific MSUs. With a base LoRA trained on the CoLA task, we adapted this LoRA for two new tasks (MNLI and MRPC) by appending an additional MSU to create two separate task-specific LoRAs. Throughout the training process for the new tasks, only the newly introduced MSU for each task was trainable. In this way, the only difference between the LoRAs for MNLI and MRPC was the unique MSU added for each, which encapsulated distinct semantic information tailored to each task. This setup was designed to create two task-specific LoRAs that differed only in one MSU, allowing us to observe parameter interference when merging these task-specific MSUs. The results, depicted in Tab.2.3, demonstrated that averaging the task-specific MSUs from the two LoRAs significantly reduced performance on each task. In contrast, maintaining these task-specific MSUs through concatenation preserved the capabilities specific to each original task. This suggests that ideal merging should maintain task-specific MSUs during LoRA merging to prevent knowledge conflict and effectively resolve parameter interference.
3 Methodology
3.1 LoRA-LEGO framework
Figure 3: The LoRA-LEGO framework merges candidate LoRAs in a manner akin to playing with LEGO by: a) first disassembling LoRAs into multiple MSUs and grouping them into an MSU pool; b) performing MSU clustering to merge similar MSUs; c) reconstructing the merged LoRA from the centroid MSUs to form a cohesive LoRA.
Based on the motivation that MSUs as the building blocks of LoRA, we can disassemble and reassemble LoRA like playing with LEGO. Here, we propose a flexible and effective method called LoRA-LEGO as shown in Fig.3. This framework is structured around three main procedures: MSU Grouping, MSU Clustering, and LoRA Reconstruction. These steps collectively facilitate the integration of diverse MSUs into a cohesive LoRA, alleviating the parameter interference while LoRA merging.
MSU Grouping.
The initial stage of merging $p$ LoRAs begins by disassembling each LoRA into various MSUs and grouping all the MSUs from each LoRA together. Let $\{{\bm{A}}_{i},{\bm{B}}_{i}\}_{i=1}^{p}$ represent the LoRA parameters of a layer with rank $r_{i}$. Each LoRA module ${{\bm{A}}_{j},{\bm{B}}_{j}}$ contains $r_{j}$ MSUs, denoted by $\{{\bm{s}}_{j1},{\bm{s}}_{j2},\ldots,{\bm{s}}_{jr_{j}}\}$, where ${\bm{s}}_{jl}=[{\bm{A}}_{jl},{\bm{B}}_{lj}^{T}]$. The MSU pool $\Phi$, which includes MSUs from all the LoRAs to be merged, is constructed as $\Phi=\bigcup_{j=1}^{k}\{{\bm{s}}_{j1},{\bm{s}}_{j2},\ldots,{\bm{s}}_{jr_{j}}\}$.
MSU Clustering.
After grouping the MSUs from different LoRAs, the next step involves regrouping these MSUs into clusters based on their similarities. With the MSU pool $\Phi$, we employed K-means Kanungo et al. (2002) to partition these MSUs into $k$ clusters $\{{\mathbb{C}}_{1},{\mathbb{C}}_{2},\ldots,{\mathbb{C}}_{k}\}$ in which each MSU is assigned to the cluster closest to it. This process is described by the following optimization problem:
$\underset{{\mathbb{C}}}{\text{minimize}}\quad\sum_{i=1}^{k}\sum_{{\bm{s}}\in{\mathbb{C}}_{i}}\|{\bm{s}}-{\bm{\mu}}_{i}\|^{2},$ (2)
where ${\bm{\mu}}_{i}$ is the centroid of cluster ${\mathbb{C}}_{i}$.
LoRA Reconstruction.
Following the MSU clustering, we rearrange the MSUs into $k$ clusters based on their similarity. The centroids of these clusters, denoted by ${\bm{\mu}}_{1},{\bm{\mu}}_{2},\ldots,{\bm{\mu}}_{k}$, are calculated as the average of the MSUs within each cluster. These centroids represent aggregated parameters across the MSUs, encapsulating the generalized semantic information most representative of each cluster. Aggregating within each cluster minimizes information loss compared to directly merging different LoRAs, as the MSUs within a cluster are more similar to each other.
Using these $k$ centroids, we can reconstruct a new LoRA module. Each centroid ${\bm{\mu}}_{i}$ contributes to a single rank in the merged model, thus the new LoRA model has a rank $k$, where $1\leq k\leq\sum_{j=1}^{p}r_{j}$. The new merged LoRA model is formed by constructing new projection matrices ${\bm{A}}^{\prime}$ and ${\bm{B}}^{\prime}$ from the centroids:
${\bm{A}}^{\prime}=\begin{bmatrix}{\bm{a}}_{1}\\ {\bm{a}}_{2}\\ \ldots\\ {\bm{a}}_{k}\end{bmatrix},\quad{\bm{B}}^{\prime}=\begin{bmatrix}{\bm{b}}_{1}^{T}&{\bm{b}}_{2}^{T}&\ldots&{\bm{b}}_{k}^{T}\end{bmatrix},$ (3)
where ${\bm{a}}_{i}$ and ${\bm{b}}_{i}$ are extracted from each centroid ${\bm{\mu}}_{i}=[{\bm{a}}_{i},{\bm{b}}_{i}]$ as per the MSU definition. The reconstructed LoRA module $\{{\bm{A}}^{\prime},{\bm{B}}^{\prime}\}$ addresses parameter interference by aligning MSUs based on their similarity before merging, achieving a flexible rank that encapsulates comprehensive knowledge across various tasks. An interesting point is that our method sits between model merging, which fuses multiple identical models into the same model, and model ensemble, which takes the average of outputs from different modules, achieving a balance between performance and computational efficiency. We provide a detailed discussion of how our method relates to model merging and model ensemble in the Appendix B.
3.2 Optimal scale of Merged LoRA
Given that the rank of the merged LoRA from LoRA-LEGO can range from $1$ to $\sum_{j=1}^{p}r_{j}$, the scale of LoRA’s output could vary significantly, thereby impacting the performance. We identified two key factors that determine the scale of the output.
Figure 4: Comparison of cluster center norm to average norm within the cluster.
Norm Decay After LoRA Merging.
As shown in Fig.4, we examine the norms of the parameters after merging (i.e., the centroids of each cluster) compared to the average norms of the parameters within each cluster before merging. We observed that after merging, the parameter norms significantly decrease, potentially affecting the output scale of the LoRA module, since the parameter norm influences the magnitude of the output. This phenomenon can be explained by the triangle inequality Klement et al. (2013), which states that for any vectors ${\bm{s}}_{i}$, $\left\|\sum_{i=1}^{p}{\bm{s}}_{i}\right\|\leq\sum_{i=1}^{p}\|{\bm{s}}_{i}\|$. When computing the centroid ${\bm{\mu}}=\frac{1}{p}\sum_{i=1}^{p}{\bm{s}}_{i}$, its norm satisfies:
$\|{\bm{\mu}}\|=\left\|\frac{1}{p}\sum_{i=1}^{p}{\bm{s}}_{i}\right\|\leq\frac{1}{p}\sum_{i=1}^{p}\|{\bm{s}}_{i}\|.$
Therefore, the norm of the centroid is less than or equal to the average of the norms of the original vectors, explaining the observed norm decay after merging. The more diverse the vectors within a cluster, the more pronounced this reduction in norm will be. To compensate for the reduced norm after merging, we perform parameter reweighting by scaling the centroid to match the average norm of the cluster: ${\bm{\mu}}^{\prime}=\frac{\frac{1}{p}\sum_{i=1}^{p}\|{\bm{s}}_{i}\|}{\|{\bm{\mu}}\|}{\bm{\mu}}$. In our implementation, we use the infinity norm for reweighting to ensure stability and robustness in the results.
Variance Expansion with Increased LoRA Rank.
Another factor influencing the scale of the LoRA output is the rank of the merged LoRA. We conducted experiments to investigate how the rank of the LoRA affects the output scale by merging seven LoRAs with rank $r=8$ and varying the rank $k$ of the merged LoRA (which corresponds to the clusters number in LoRA-LEGO). The frequency histograms of outputs from the first layer of the merged LoRA at various ranks, as shown in Fig.5, indicate that LoRA outputs approximate a normal distribution centered at zero. We observed that as the rank $k$ increases, the variance of the output also increases. To normalize the output variance, similar to the normalization in the self-attention mechanisms Vaswani (2017), we perform output reweighting for the merged LoRA by the factor $\frac{\sqrt{r}}{\sqrt{k}}$. The following theorem ensures that this rescaling maintains a consistent variance in the LoRA output.
Figure 5: Expansion of variance with increasing rank in merged LoRAs.
Theorem 3.1.
Let ${\bm{A}}_{1}\in\mathbb{R}^{p\times r}$ and ${\bm{B}}_{1}\in\mathbb{R}^{r\times p}$, and ${\bm{A}}_{2}\in\mathbb{R}^{p\times k}$ and ${\bm{B}}_{2}\in\mathbb{R}^{k\times p}$, where all elements of these matrices are independently and identically distributed according to the standard normal distribution $\mathcal{N}(0,1)$. Then, after scaling the product ${\bm{A}}_{2}{\bm{B}}_{2}$ by the factor $\sqrt{r}/\sqrt{k}$, the variances of the entries of ${\bm{A}}_{1}{\bm{B}}_{1}$ and the scaled ${\bm{A}}_{2}{\bm{B}}_{2}$ are equal:
$\mathrm{Var}\left({\bm{A}}_{1}{\bm{B}}_{1}\right)=\mathrm{Var}\left(\frac{\sqrt{r}}{\sqrt{k}}{\bm{A}}_{2}{\bm{B}}_{2}\right).$
The proof of Theorem 3.1 is detailed in the Appendix D. Overall, to ensure that the LoRA output is correctly scaled, we employ two scaling strategies. First, we reweight the parameters to match the average norms of the parameters within each cluster. Second, we rescale the output of each LoRA to the rank of the merged LoRA, thereby maintaining variance consistency with the original LoRA. These dual scaling strategies enable LoRA-LEGO to deliver enhanced and more robust performance.
4 Experiments
Given that LoRA merging is essential for many scenarios, we have opted for two settings: Multi-task Learning Tang et al. (2024) and Mixed-task Settings Zhao et al. (2024b). In these settings, we compared various LoRA composition methods to assess the performance of the proposed LoRA-LEGO approach. We selected Llama2-{7b,13b} as the base model and trained LoRA for each task with hyperparameters $r=6$ and $\alpha=12$. The evaluation frameworks for multi-task Learning and mixed-task settings are detailed in the subsequent sections, where we provide a comprehensive analysis.
4.1 Multi-task Learning
Experiment Setting.
Multi-task Learning aims to merge individually trained LoRAs into a unified model while preserving the performance of each constituent LoRA. Drawing from previous research Tang et al. (2024); Yadav et al. (2024b); Ilharco et al. (2022), we merged seven LoRA models, each fine-tuned on Llama2-7b,13b, for in-domain tasks including Cola, Mnli, MRPR, QNLI, GLUE-QQP, RTE, and SST2. We then assessed the performance of the merged LoRA on these in-domain tasks as well as on two additional out-of-domain tasks, SNLI and WNLI, to evaluate its adaptability and generalization capabilities.
Baseline Methods.
We compared the proposed method with four post-hoc training-free LoRA composition methods, including (1) Weight Averaging, (2) Ensemble, (3) Task Arithmetic, and (4) Ties-Merging. The details of these LoRA composition methods can be found in the Appendix C.
Table 3: Multi-task performance when merging Llama2-{7b,13b} (LoRA fine-tuned) models on seven seen tasks and two unseen tasks.
IID Tasks OOD Tasks Average
Method CoLA MNLI MRPC QNLI QQP RTE SST2 SNLI WNLI
w/ Llama2-7b
Task LoRA 61.63 77.46 68.00 77.25 75.83 52.22 75.74
Weight Average 54.42 36.09 68.00 44.41 51.72 48.15 42.99 31.64 47.14 47.17
Ensemble 55.67 45.89 59.25 59.84 67.38 68.89 66.44 36.73 51.43 56.84
Task Arithmetic 55.48 42.15 54.25 58.94 66.43 67.78 59.54 34.08 54.29 54.77
Ties-Mering 48.65 48.81 55.50 61.79 66.75 62.59 70.69 48.45 61.43 58.30
LoRA-LEGO 55.48 55.73 66.00 62.29 71.07 71.85 73.22 51.36 52.86 62.21
w/ Llama2-13b
Task LoRA 69.04 88.23 89.25 82.33 86.29 80.74 76.44
Weight Average 45.48 46.32 67.75 46.68 47.50 62.96 46.78 42.42 42.86 49.86
Ensemble 62.50 64.64 74.75 71.81 81.35 79.26 75.52 54.32 60.00 69.35
Task Arithmetic 63.17 64.41 74.50 71.59 80.84 78.15 75.86 54.16 58.57 69.03
Ties-Mering 58.56 64.71 78.75 74.27 80.71 76.67 75.40 56.02 61.43 69.61
LoRA-LEGO 59.42 65.40 75.50 72.29 82.51 78.52 75.98 58.54 64.29 70.27
Figure 6: LoRA pruning performance over seven datasets.
Main Results.
As shown in Tab. 3, our proposed LoRA-LEGO method significantly outperforms the baseline methods on both IID and OOD tasks. Specifically, the Weight Averaging method suffers from significant performance degradation due to parameter interference during LoRA merging. The Ensemble method encounters issues with parameter redundancy, leading to suboptimal performance and slower inference speeds. Model merging methods such as Task Arithmetic and Ties-Merging perform element-wise fusion and fail to adequately address parameter interference in LoRA, resulting in suboptimal performance during the merging process. In contrast, our proposed LoRA-LEGO effectively alleviates parameter misalignment and knowledge conflict through flexible MSU clustering, thereby achieving superior performance compared to other methods. In Appendix E, we demonstrate that the proposed LoRA-LEGO approach can effectively merge heterogeneous LoRAs, exceeding the capabilities of previous model merging methods.
Figure 7: Ablation on scaling strategies.
Performance on LoRA Pruning.
Our method also functions as a LoRA parameter pruning approach. For a single LoRA with rank $r$, LoRA-LEGO allows for selecting $k<r$, effectively reducing the rank to $k$ and pruning the model. As illustrated in Fig. 6, we evaluate the performance of a single LoRA model after retaining various proportions of its parameters. LoRA-LEGO efficiently compresses model parameters: retaining just 33% of the parameters preserves 79% of the original model’s capabilities while keeping 50% maintains 92% of the performance. This offers new insights into strategies for compressing model parameters, especially those of LoRA.
Ablation of Scaling Strategies.
We evaluate the effectiveness of two scaling strategies for the merged LoRA by varying the number of clusters for LoRA-LEGO, noting that the cluster number corresponds to the rank of the merged LoRA. As illustrated in Fig.7, the original computation of LoRA experiences significant performance degradation with increasing rank of the merged LoRA, primarily due to the expansion of variance associated with the increased rank. Additionally, when the rank of the merged LoRA is relatively low, its performance does not reach its optimum due to the degradation of parameter norms. We also present the performance of each scaling strategy and their combination. Applying parameter reweighing can significantly enhance the performance of the merged LoRA; specifically, the performance of a merged LoRA at rank $1r$ improves by 5%. However, performance degradation still occurs at higher LoRA ranks due to variance expansion. Stabilizing the variance by output reweighting significantly increases performance when the rank is high, although it remains suboptimal due to the decrease of parameter norms. Combining these two scaling strategies yields the best results, demonstrating stable and improved performance across varying ranks of the merged LoRA. After these two scaling strategies are applied, the performance of LoRA-LEGO tends to stabilize; therefore, we use $k=2r$ as the default setting.
Merging Different Number of Tasks.
We investigated the average performance of the model when merging LoRAs with different numbers of tasks. To better assess the influence of task quantity on our method, we normalized the performance of each task by dividing it by the performance of its respective single-task LoRA and then calculated the mean of these normalized scores. From Fig.8, it is evident that as the number of merging tasks increases, there is a general decline in the performance of the merged LoRAs. Specifically, direct averaging experiences a steep performance drop due to parameter interference. The Ensemble method also sees a decrease in performance, attributed to parameter redundancy and inconsistencies in the output space. Ties-merging, failing to resolve parameter interference and reliant on hyperparameter selection fully, does not reach optimal performance. LoRA-LEGO, which flexibly addresses parameter interference, experiences a lesser decline in performance with an increasing number of tasks, thereby outperforming the baseline model.
Figure 8: Average performance varying the number of merged tasks.
Table 4: The average performance of each task cluster. The performance of perfectly selected corresponding LoRA for each sample is colored in gray. We have bolded the best performance of each task and underlined the best performance in the “OOD” setting.
Task Perfect Selection LoRA-LEGO Selection Weight Average Ensemble Tie-Merging
IID OOD IID OOD IID OOD IID OOD IID OOD
w/ Llama2-7b
Struct to Text Rouge-1 59.1 49.6 50.5 56.8 45.2 44.5 41.0 51.2 45.3 45.4 49.4
Struct to Text Rouge-2 36.1 25.7 26.6 33.6 23.2 22.6 20.2 26.3 22.9 23.9 27.4
Struct to Text Rouge-l 48.6 39.5 39.9 46.4 35.3 34.5 31.7 41.0 35.5 36.0 39.3
Translation BLEU 13.1 12.9 12.4 12.8 12.0 12.2 12.3 12.8 12.2 14.0 13.9
COMMONSENSE 62.5 60.0 60.5 55.5 46.0 51.0 48.0 61.5 50.0 55.0 59.5
SENTIMENT 90.0 90.0 91.5 89.5 89.0 79.0 78.5 89.5 90.5 82.0 81.5
READING Comp. 67.3 54.3 55.7 51.7 40.3 47.3 45.0 51.3 47.3 46.3 56.3
CLOSE-BOOK QA 45.0 47.0 48.5 40.0 43.0 41.0 37.5 45.0 48.5 48.0 53.5
COREFERENCE 52.0 62.0 60.0 50.0 46.0 47.0 53.0 63.0 49.0 32.0 47.0
READ. COOMP. W/ COM 69.0 66.0 65.0 69.0 30.0 35.0 19.0 46.0 40.0 37.0 64.0
PARAPHRASE 65.5 58.0 60.0 58.0 45.5 45.5 44.0 56.5 45.5 18.0 38.5
NLI 72.3 71.3 66.4 70.0 60.6 51.4 53.8 67.9 64.3 65.6 49.4
Overall 55.4 51.4 51.0 51.2 43.0 41.6 40.2 49.8 45.6 43.2 45.8
w/ Llama2-13b
$\text{Struct to Text}_{Rouge-1}$ 61.0 54.2 46.0 58.0 44.6 48.2 45.1 52.9 46.9 50.8 50.9
$\text{Struct to Text}_{Rouge-2}$ 37.7 29.3 24.0 34.9 22.8 26.0 23.5 29.1 24.6 26.2 26.1
$\text{Struct to Text}_{Rouge-l}$ 50.5 43.9 36.4 47.6 34.8 38.4 35.9 42.9 36.9 41.0 40.9
$\text{Translation}_{BLEU}$ 12.9 14.7 14.5 12.9 12.7 14.6 14.1 14.6 14.1 11.2 11.3
COMMONSENSE 69.5 69.0 68.5 59.0 47.5 61.0 56.0 64.0 60.5 58.0 57.5
SENTIMENT 90.0 91.0 90.0 90.5 91.0 87.0 83.5 91.5 91.5 91.5 91.5
READING Comp. 76.0 62.7 53.0 60.3 48.0 56.7 49.3 60.3 51.3 54.3 54.3
CLOSE-BOOK QA 64.0 63.0 58.0 60.0 53.0 62.0 58.0 63.0 61.0 41.5 42.0
COREFERENCE 74.0 77.0 62.0 75.0 65.0 55.0 59.0 76.0 64.0 63.0 63.0
READ. COOMP. W/ COM 82.0 76.0 54.0 80.0 33.0 57.0 49.0 78.0 58.0 65.0 66.0
PARAPHRASE 77.5 67.5 58.5 68.0 52.5 55.5 45.5 71.0 55.5 61.0 62.5
NLI 82.4 78.9 76.3 78.9 70.2 69.8 66.4 78.1 75.7 65.7 65.7
Overall 62.4 58.2 52.8 57.8 47.7 51.6 47.8 57.6 52.3 50.1 50.3
4.2 Mixed-task Evaluation
Evaluation Setting.
Recent studies Zhao et al. (2024b) have proposed the creation of a LoRA pool from which relevant LoRAs are retrieved for each input to facilitate LoRA composition. We adopt the same setting and construct a LoRA pool for 48 tasks from flan-v2, grouped into 10 task clusters. The evaluation set is constructed by randomly choosing 50 samples from each test set. These samples are then mixed and shuffled to form a unified dataset comprising 2400 data points.
Adopting the LoraRetriever approach Zhao et al. (2024b), we initially retrieve the top-3 LoRAs based on the sentence embedding similarities between each input sample and a corresponding few-shot samples. Following this, we engage in LoRA composition and evaluate various strategies. This analysis underscores the versatility and superior performance of LoRA-LEGO in handling more complex scenarios.
Baseline Methods.
For all methods, we employ a consistent evaluation pipeline. For each instance in the evaluation set, we initially retrieve the top-3 LoRA, followed by the composition of LoRA. We compared the following LoRA composition methods: (1) Weighted Average, (2) Ensemble, (3) Selection (using the top-1 retrieved LoRA), and (4) Ties-Merging.
Main Results.
Previous research Zhao et al. (2024b) has shown that using a retriever to identify LoRA tasks tailored to various inputs is more efficient and effective in personalized service settings. Consequently, we concentrate on how multiple LoRAs can be integrated effectively through LoRA merging after retrieving the top-k LoRAs for each input. We assess the performance of LoRA composition methods in both IID and OOD contexts. “IID” performance refers to scenarios where all LoRAs are accessible to the retriever. “OOD” performance, however, involves masking the LoRA associated with the specific task of each test sample during retrieval, preventing any sample from accessing its ideal LoRA. This approach allows us to evaluate the cross-task generalization capabilities of the LoRA composition methods. Tab.4 demonstrates that LoRA-LEGO surpasses other composition methods in both IID and OOD scenarios by fully eliminating parameter interference. In contrast, baseline LoRA composition methods experience performance degradation due to their inability to completely mitigate parameter interference. Specifically, in IID scenarios, the Selection method excels because the Retriever can choose the most appropriate LoRA from closely related tasks for inference. Building on this, LoRA-LEGO further enhances performance by leveraging the transfer capabilities between different tasks, thereby achieving better results. For OOD scenarios, both Ties-Merging and Ensemble show good performance by harnessing knowledge from a wide array of relevant tasks to tackle OOD tasks. LoRA-LEGO, however, outperforms these methods by effectively addressing parameter interference, allowing for a more comprehensive utilization of diverse LoRA capabilities and achieving superior results in OOD setting.
5 Related Work
Model Merging.
Many works have discussed how to obtain a comprehensive model through model merging from various perspectives. Some works discuss how to find a set of low-loss paths in the parameter space for model parameter interpolation from the perspective of linear mode connectivity Ainsworth et al. (2022); Entezari et al. (2021). From a similar perspective, we further utilized properties of MSUs, mploying clustering algorithms to provide a flexible solution for enhancing the parameter connectivity during LoRA merging. Additionally, many works attempt to coordinate models trained in a decentralized and separated manner through model merging, utilizing their knowledge transfer capabilities to obtain a model with comprehensive abilities Tang et al. (2024); Don-Yehiya et al. (2022); Yadav et al. (2024b); Matena & Raffel (2022); Jin et al. (2022); Yang et al. (2023a). Recently, with the rise of large language models, more and more works have focused on how to use model aggregation, especially the aggregation of LoRA Chronopoulou et al. (2023); Huang et al. (2023); Zhao et al. (2024b); Wang et al. (2024), to strategically utilize models adapted to multiple domains. These efforts often overlook the parameter interference that occurs during LoRA merging, and some of them require extensive additional training or adaptation. This leads to suboptimal performance in such scenarios or restricts their applicability.
Application of LoRA Merging.
LoRA merging can be applied in various scenarios. For instance, in multi-task learning Tang et al. (2024); Don-Yehiya et al. (2022), models adapt to different domains in a decentralized manner using LoRA, subsequently acquiring multi-task capabilities through merging. In mixed-task scenarios Zhao et al. (2024b; a), LoRAs from diverse domain tasks are uploaded to a centralized service platform, where the service retrieves and composes LoRAs to deliver personalized services based on downstream requests. In federated learning Chen et al. (2023); Zhang et al. (2024), edge devices train LoRAs on private data and upload them to a central server for merging and distribution, enabling iterative optimization through this process. During the alignment phase, RLHF training is conducted for various preferences to obtain multiple LoRAs that meet different requirements. This is followed by providing personalized alignment models through parameter interpolation Jang et al. (2023).
6 Conclusion
In this paper, we address the critical challenge of merging multiple LoRAs, each tailored for distinct tasks, into a unified and comprehensive LoRA. We identify parameter interference as a primary obstacle in merging, with parameter misalignment and knowledge conflict being significant contributors. Our exploration of LoRA’s properties reveals several key insights: (1) Each rank within a LoRA operates independently and represents a minimal semantic unit (MSU); (2) MSUs within each LoRA exhibit permutation invariance; (3) MSUs can be concatenated to form a comprehensive LoRA. Building on these insights, we propose LoRA-LEGO, a methodology that aggregates MSUs from all target LoRAs, performs clustering, and uses the centroid of each cluster to create a merged LoRA. Our extensive experimental results validate the effectiveness of the LoRA-LEGO approach.
Potential future work includes exploring alternative distance metrics for LoRA-LEGO, such as optimal transport, to provide a more detailed characterization of parameter similarities beyond the typical Euclidean distance. Additionally, further modularization of LoRA could enhance various applications, particularly in federated learning where strategies to minimize communication overhead and expedite model convergence through sharing and aggregating MSUs might be explored. We believe these advancements can significantly benefit a wide range of fields and applications.
Appendix A Difference between LoAR Merging Setting and Model Merging Setting
Previous work on model merging primarily focused on integrating separately trained models to form a comprehensive system. These methods typically involve reloading LoRA parameters into the original model before merging, which introduces additional overhead by necessitating the reconstruction of a corresponding LLM for each LoRA. In many cases, the goal of LoRA merging is to create a new LoRA that consolidates the capabilities of all involved LoRAs for simplified task-specific usage. In contrast, the LoRA merging approach presented in this paper bypasses this step; it directly merges the LoRA parameters to generate a unified LoRA with enhanced capabilities.
Appendix B Connection with Vanilla LoRA Composition Methods
The vanilla LoRA composition comprises two training-free methods: the Ensemble of LoRAs and the Fusion of LoRAs. The Ensemble of LoRAs strategy involves the aggregation of the outputs of each submodule within the assembled LoRAs. Let us denote ${\mathbb{A}}=\{{\bm{A}}_{1},{\bm{A}}_{2},\ldots,{\bm{A}}_{n}\}$ and ${\mathbb{B}}=\{{\bm{B}}_{1},{\bm{B}}_{2},\ldots,{\bm{B}}_{n}\}$ as the sets representing submodules within $n$ LoRAs. For an input ${\bm{x}}_{i}$, the output derived from the Ensemble of LoRAs can be expressed as ${\bm{x}}_{i}^{\prime}=\frac{1}{n}\sum_{j=1}^{n}{\bm{B}}_{j}{\bm{A}}_{j}{\bm{x}}_{i}$, where ${\bm{x}}_{i}^{\prime}$ denotes the output.
In contrast to the Ensemble method, which combines the output of different LoRAs, model merging presents an alternative composition strategy. Let the parameters of each LoRA, $\phi_{i}$, be denoted by $\Theta_{i}$. A typical strategy involves averaging these parameters, represented as $\Theta_{\text{fusion}}=\frac{1}{k}\sum_{j=1}^{k}\Theta_{j}$. This formulation allows the fused parameter to function akin to a single LoRA.
Our proposed LoRA-LEGO method offers refined control over model complexity and performance, facilitating an optimal balance between semantic integration and computational efficiency. Unlike Simple Fusion, which lacks alignment and may introduce inconsistencies across merged LoRAs, our approach aligns and selectively fuses MSUs based on their semantic closeness. This selective fusion significantly reduces redundancy, a prevalent issue in the Ensemble method where all MSUs are treated uniformly, potentially leading to an excessive parameter count that can degrade performance. Our clustering-based strategy addresses these challenges by effectively condensing the most relevant semantic features into fewer clusters, thereby enhancing both the model’s performance and its inference efficiency.
Appendix C Details of Baseline Methods
We compare our method with the following baseline:
  1. 1.
    Weight Averaging. This approach averages the parameters across different instances of LoRA, resulting in a new composite LoRA defined as ${\bm{A}}^{\prime}=\frac{1}{n}\sum_{i=1}^{n}{\bm{A}}_{i}$ and ${\bm{B}}^{\prime}=\frac{1}{n}\sum_{i=1}^{n}{\bm{B}}_{i}$, where ${\bm{A}}_{i}$ and $B_{i}$ represent the parameters from the $i$-th instance of the original LoRA models, and $n$ is the number of models being averaged.
  2. 2.
    Ensemble. This method averages the outputs from each LoRA, simultaneously activating multiple LoRAs to compose a combined output. The specific calculation for the mixed output is defined as ${\bm{x}}^{\prime}=\frac{1}{n}\sum_{i=1}^{n}{\bm{B}}_{j}{\bm{A}}_{j}{\bm{x}}_{i}$.
  3. 3.
    Task Arithmetic. This method is akin to weight averaging, but it differentiates by using weights derived from a hyper-parameter search to merge models. The calculations for this composite are ${\bm{A}}^{\prime}=p\sum_{i=1}^{n}{\bm{A}}_{i}$ and ${\bm{B}}^{\prime}=p\sum_{i=1}^{n}{\bm{B}}_{i}$, where $p$ represents the hyper-parameter that scales the contributions of each model.
  4. 4.
    Ties-Merging. This method aims to resolve element-wise parameter interference by initially trimming the redundant parameters, retaining only the top-kof values based on their magnitude. It then selects the sign vector for the merged model and finally performs a disjoint mean operation. Ties-Merging posits that the primary source of parameter interference arises from inconsistencies in the values of merged parameters, while potentially overlooking issues related to misalignment and knowledge conflict.
Appendix D Optimal Scale of Merged LoRA
Theorem D.1.
Let ${\bm{A}}_{1}\in\mathbb{R}^{p\times r}$ and ${\bm{B}}_{1}\in\mathbb{R}^{r\times p}$, and ${\bm{A}}_{2}\in\mathbb{R}^{p\times k}$ and ${\bm{B}}_{2}\in\mathbb{R}^{k\times p}$, where all elements of these matrices are independently and identically distributed according to the standard normal distribution $\mathcal{N}(0,1)$. Then, after scaling the product ${\bm{A}}_{2}{\bm{B}}_{2}$ by the factor $\sqrt{r}/\sqrt{k}$, the variances of the entries of ${\bm{A}}_{1}{\bm{B}}_{1}$ and the scaled ${\bm{A}}_{2}{\bm{B}}_{2}$ are equal:
$\mathrm{Var}\left({\bm{A}}_{1}{\bm{B}}_{1}\right)=\mathrm{Var}\left(\frac{\sqrt{r}}{\sqrt{k}}{\bm{A}}_{2}{\bm{B}}_{2}\right).$
Proof.
To compute the variance of the entries of the matrices ${\bm{A}}_{1}{\bm{B}}_{1}$ and $\frac{\sqrt{r}}{\sqrt{k}}{\bm{A}}_{2}{\bm{B}}_{2}$, we examine each entry individually.
For ${\bm{A}}_{1}{\bm{B}}_{1}$, each entry is calculated as:
$({\bm{A}}_{1}{\bm{B}}_{1})_{ij}=\sum_{l=1}^{r}({\bm{A}}_{1})_{il}({\bm{B}}_{1})_{lj}.$
Since $({\bm{A}}_{1})_{il}$ and $({\bm{B}}_{1})_{lj}$ are independent and follow $\mathcal{N}(0,1)$, their product has mean zero and variance one:
$\mathbb{E}\left[({\bm{A}}_{1})_{il}({\bm{B}}_{1})_{lj}\right]=0,\quad\mathrm{Var}\left(({\bm{A}}_{1})_{il}({\bm{B}}_{1})_{lj}\right)=1.$
The terms $({\bm{A}}_{1})_{il}({\bm{B}}_{1})_{lj}$ are independent for different $l$, so the variance of $({\bm{A}}_{1}{\bm{B}}_{1})_{ij}$ is:
$\mathrm{Var}\left(({\bm{A}}_{1}{\bm{B}}_{1})_{ij}\right)=\sum_{l=1}^{r}\mathrm{Var}\left(({\bm{A}}_{1})_{il}({\bm{B}}_{1})_{lj}\right)=r\times 1=r.$
Similarly, for ${\bm{A}}_{2}{\bm{B}}_{2}$, each entry is:
$({\bm{A}}_{2}{\bm{B}}_{2})_{ij}=\sum_{l=1}^{k}({\bm{A}}_{2})_{il}({\bm{B}}_{2})_{lj},$
and each term $({\bm{A}}_{2})_{il}({\bm{B}}_{2})_{lj}$ has variance one. Therefore, the variance of $({\bm{A}}_{2}{\bm{B}}_{2})_{ij}$ is:
$\mathrm{Var}\left(({\bm{A}}_{2}{\bm{B}}_{2})_{ij}\right)=\sum_{l=1}^{k}\mathrm{Var}\left(({\bm{A}}_{2})_{il}({\bm{B}}_{2})_{lj}\right)=k\times 1=k.$
After scaling ${\bm{A}}_{2}{\bm{B}}_{2}$ by $\sqrt{r}/\sqrt{k}$, the variance becomes:
$\mathrm{Var}\left(\left(\frac{\sqrt{r}}{\sqrt{k}}{\bm{A}}_{2}{\bm{B}}_{2}\right)_{ij}\right)=\left(\frac{\sqrt{r}}{\sqrt{k}}\right)^{2}\mathrm{Var}\left(({\bm{A}}_{2}{\bm{B}}_{2})_{ij}\right)=\left(\frac{r}{k}\right)\times k=r.$
Thus, the variances of the entries are equal:
$\mathrm{Var}\left({\bm{A}}_{1}{\bm{B}}_{1}\right)=\mathrm{Var}\left(\frac{\sqrt{r}}{\sqrt{k}}{\bm{A}}_{2}{\bm{B}}_{2}\right).$
Appendix E Performance on Merging Heterogeneous LoRAs
Table 5: Multi-task performance when merging heterogeneous LoRAs on seven seen tasks and two unseen tasks.
IID Tasks OOD Tasks Average
Method CoLA MNLI MRPC QNLI QQP RTE SST2 SNLI WNLI
w/ Llama2-7b
Task LoRA 61.63 77.46 68.00 82.69 75.83 77.04 77.47
Weight Average
Task Arithmetic
Ties-Mering
Ensemble 56.06 55.84 69.75 64.91 74.85 74.44 70.92 46.19 52.86 62.87
LoRA-LEGO 55.10 60.67 69.25 67.29 65.61 67.04 74.83 57.82 52.86 63.39
Another advantage of LoRA-LEGO is its ability to merge heterogeneous LoRAs, that is, LoRAs with different ranks. To experimentally verify this feature, we retrained LoRAs for the QNLI, RTE, and SST2 tasks with $r=16$ and $\alpha=32$, and merged them with LoRAs from other tasks ($r=8$, $\alpha=16$) to obtain a new LoRA. Since other model merging methods require the merged LoRAs to have the same architecture, we only compared our method with the Ensemble method. As shown in Tab.5, the results demonstrate that our method can effectively merge heterogeneous LoRAs and achieves better overall performance than the Ensemble method.

### v3

@@
 (2022); Yang et al.
-(2024) offers an alternative approach by aggregating the parameters of multiple LoRAs into a unified adapter without extra training, producing a unified LoRA with more comprehensive capabilities.
+(2024) offers an alternative approach by aggregating the parameters of multiple LoRAs into a unified adapter without extra training, producing a unified LoRA with comprehensive capabilities.
 However, these methods typically employ element-wise parameter fusion, which can neglect and disrupt the internal semantic structure within LoRA.
@@
 This paper approaches LoRA merging from a novel perspective, focusing on the fine-grained modularization of LoRA by decomposing it into independent units, which enables the flexible reconstruction of a unified LoRA with comprehensive capabilities.
-Figure 1: Further Modularization of LoRA: a) Each LoRA can be further modularized into multiple Minimal Semantic Units (MSUs), each corresponding to a column in ${\bm{A}}$ matrix and a row in matrix ${\bm{B}}$, differentiated by distinct colors.
+Figure 1: Further Modularization of LoRA: a) Each LoRA can be further modularized into multiple Minimal Semantic Units (MSUs), each corresponding to a row in ${\bm{A}}$ matrix and a column in matrix ${\bm{B}}$, differentiated by distinct colors.
 b) The MSUs within a LoRA display permutation invariance, implying that any rearrangement of the MSUs does not affect the output generated by the LoRA.
 c) Multiple LoRAs exhibit Concatenation-Summation Equivalence, indicating that the summation of outputs from various LoRAs is equivalent to the output of a singular LoRA constructed by concatenating their MSUs.
-As illustrated in Fig.1, our motivation for further modularizing LoRA stems from the following insights: a) A rank of LoRA corresponds to a column in the down-projection matrix ${\bm{A}}$ and a row in the up-projection matrix ${\bm{B}}$.
-Therefore, we consider the parameters associated with each rank as a cohesive entity and define them as Minimal Semantic Units (MSUs), which serve as the fundamental building blocks of each LoRA.
+As illustrated in Fig.1, our motivation for further modularizing LoRA stems from the following insights: a) Each rank in LoRA corresponds to a row in the down-projection matrix $\mathbf{A}$ and a column in the up-projection matrix $\mathbf{B}$.
+Since the calculations for each rank are independent, we consider the parameters associated with each rank as a cohesive entity.
+We define these entities as Minimal Semantic Units (MSUs), which serve as the fundamental building blocks of LoRA.
 b) Within each LoRA, the MSUs exhibit the property of Permutation Invariance, indicating that any permutation of MSUs within a LoRA does not affect the adapter’s output.
-c) LoRA exhibits the Concatenation-Summation Equivalence property, which states that summing the outputs from multiple LoRAs is equivalent to the output of a single higher-ranked LoRA constructed by concatenating the MSUs of these LoRAs.
+c) LoRA exhibits the Concatenation-Summation Equivalence property, which states that summing the outputs from multiple LoRAs is equivalent to the output of a single higher-ranked LoRA constructed by concatenating all the MSUs of these LoRAs.
 In this paper, we introduce a novel method called LoRA-LEGO, which is based on the insight that MSUs act as building blocks that form a LoRA and can be disassembled and reassembled like playing with LEGO.
@@
 This approach allows for targeted rank adjustments in the merged LoRA to preserve task-specific knowledge.
+We also observed that variations in parameter norms and the rank size of the merged LoRA affect the output scale.
+To address this, we implement a dual reweighting strategy that adjusts both the parameters and the outputs, ensuring optimal scaling for the merged LoRA.
 We empirically validate the effectiveness of the proposed LoRA-LEGO in both multi-task Tang et al.
@@
 Let ${\bm{A}}$ and ${\bm{B}}$ be matrices in a LoRA module.
-For each index $i$, define the minimum semantic unit of LoRA as the combined vector ${\bm{s}}_{i}=[{\bm{a}}_{i},{\bm{b}}_{i}]$, where ${\bm{a}}_{i}$ is the $i$-th column of ${\bm{A}}$ and ${\bm{b}}_{i}$ is the $i$-th column of ${\bm{B}}^{T}$ (i.e., the transpose of the $i$-th row of ${\bm{B}}$).
-In this context, each column of the down-projection matrix ${\bm{A}}$ and its corresponding row in the up-projection matrix ${\bm{B}}$ are treated as a cohesive unit, defined as a Minimum Semantic Unit (MSU).
+For each index $i$, define the minimum semantic unit of LoRA as the combined vector ${\bm{s}}_{i}=[{\bm{a}}_{i},{\bm{b}}_{i}]$, where ${\bm{a}}_{i}$ is the $i$-th row of ${\bm{A}}$ and ${\bm{b}}_{i}$ is the $i$-th row of ${\bm{B}}^{T}$ (i.e., the transpose of the $i$-th column of ${\bm{B}}$).
+In this context, each row of the down-projection matrix ${\bm{A}}$ and its corresponding column in the up-projection matrix ${\bm{B}}$ are treated as a cohesive unit, defined as a Minimum Semantic Unit (MSU).
 Each MSU contributes to a rank of the LoRA, encapsulating a distinct semantic fragment of the LoRA’s capacity.
@@
 This setup was designed to create two task-specific LoRAs that differed only in one MSU, allowing us to observe parameter interference when merging these task-specific MSUs.
-The results, depicted in Tab.2.3, demonstrated that averaging the task-specific MSUs from the two LoRAs significantly reduced performance on each task.
+The results, depicted in Tab.2, demonstrated that averaging the task-specific MSUs from the two LoRAs significantly reduced performance on each task.
 In contrast, maintaining these task-specific MSUs through concatenation preserved the capabilities specific to each original task.
@@
 Let $\{{\bm{A}}_{i},{\bm{B}}_{i}\}_{i=1}^{p}$ represent the LoRA parameters of a layer with rank $r_{i}$.
-Each LoRA module ${{\bm{A}}_{j},{\bm{B}}_{j}}$ contains $r_{j}$ MSUs, denoted by $\{{\bm{s}}_{j1},{\bm{s}}_{j2},\ldots,{\bm{s}}_{jr_{j}}\}$, where ${\bm{s}}_{jl}=[{\bm{A}}_{jl},{\bm{B}}_{lj}^{T}]$.
+Each LoRA module ${{\bm{A}}_{j},{\bm{B}}_{j}}$ contains $r_{j}$ MSUs, denoted by $\{{\bm{s}}_{j1},{\bm{s}}_{j2},\ldots,{\bm{s}}_{jr_{j}}\}$, where ${\bm{s}}_{jl}=[{\bm{a}}_{jl},{\bm{b}}_{jl}]$ with ${\bm{a}}_{jl}={\bm{A}}_{j}[:,l]$ and ${\bm{b}}_{jl}={\bm{B}}_{j}[l,:]^{T}$.
 The MSU pool $\Phi$, which includes MSUs from all the LoRAs to be merged, is constructed as $\Phi=\bigcup_{j=1}^{k}\{{\bm{s}}_{j1},{\bm{s}}_{j2},\ldots,{\bm{s}}_{jr_{j}}\}$.
@@
 The reconstructed LoRA module $\{{\bm{A}}^{\prime},{\bm{B}}^{\prime}\}$ addresses parameter interference by aligning MSUs based on their similarity before merging, achieving a flexible rank that encapsulates comprehensive knowledge across various tasks.
-An interesting point is that our method sits between model merging, which fuses multiple identical models into the same model, and model ensemble, which takes the average of outputs from different modules, achieving a balance between performance and computational efficiency.
+An interesting point is that our method sits between model merging, which fuses multiple identical models into a singular model, and model ensemble, which takes the average of outputs from different modules, achieving a balance between performance and computational efficiency.
 We provide a detailed discussion of how our method relates to model merging and model ensemble in the Appendix B.
@@
 First, we reweight the parameters to match the average norms of the parameters within each cluster.
-Second, we rescale the output of each LoRA to the rank of the merged LoRA, thereby maintaining variance consistency with the original LoRA.
+Second, we rescale the output of the merged LoRA for maintaining variance consistency with the original LoRA.
 These dual scaling strategies enable LoRA-LEGO to deliver enhanced and more robust performance.
@@
 We also present the performance of each scaling strategy and their combination.
-Applying parameter reweighing can significantly enhance the performance of the merged LoRA; specifically, the performance of a merged LoRA at rank $1r$ improves by 5%.
-However, performance degradation still occurs at higher LoRA ranks due to variance expansion.
+Applying parameter reweighting can significantly enhance the performance of the merged LoRA when the rank is relatively low; specifically, the performance of a merged LoRA at rank $1r$ improves by 5%.
+However, as the rank increases, eliminating norm decay more severely exposes variance expansion because norm decay can alleviate this phenomenon, leading to greater performance degradation.
 Stabilizing the variance by output reweighting significantly increases performance when the rank is high, although it remains suboptimal due to the decrease of parameter norms.
@@
 Adopting the LoraRetriever approach Zhao et al.
-(2024b), we initially retrieve the top-3 LoRAs based on the sentence embedding similarities between each input sample and a corresponding few-shot samples.
+(2024b), we initially retrieve the top-3 LoRAs based on the sentence embedding similarities between each input sample and LoRA’s few-shot samples.
 Following this, we engage in LoRA composition and evaluate various strategies.
@@
 For each instance in the evaluation set, we initially retrieve the top-3 LoRA, followed by the composition of LoRA.
-We compared the following LoRA composition methods: (1) Weighted Average, (2) Ensemble, (3) Selection (using the top-1 retrieved LoRA), and (4) Ties-Merging.
+We compared the following LoRA composition methods: (1) Weight Average, (2) Ensemble, (3) Selection (using the top-1 retrieved LoRA), and (4) Ties-Merging.
 Main Results.
@@
 (2021).
-From a similar perspective, we further utilized properties of MSUs, mploying clustering algorithms to provide a flexible solution for enhancing the parameter connectivity during LoRA merging.
+From a similar perspective, we further utilized properties of MSUs, employing clustering algorithms to provide a flexible solution for enhancing the parameter connectivity during LoRA merging.
 Additionally, many works attempt to coordinate models trained in a decentralized and separated manner through model merging, utilizing their knowledge transfer capabilities to obtain a model with comprehensive abilities Tang et al.
@@
 (2024), edge devices train LoRAs on private data and upload them to a central server for merging and distribution, enabling iterative optimization through this process.
-During the alignment phase, RLHF training is conducted for various preferences to obtain multiple LoRAs that meet different requirements.
-This is followed by providing personalized alignment models through parameter interpolation Jang et al.
+During the alignment phase, Reinforcement Learning from Human Feedback (RLHF) training is conducted to obtain multiple LoRA models that meet different requirements based on various preferences.
+Subsequently, personalized alignment models can be provided through parameter interpolation, as discussed in  Jang et al.
 (2023).
@@
 Our extensive experimental results validate the effectiveness of the LoRA-LEGO approach.
-Potential future work includes exploring alternative distance metrics for LoRA-LEGO, such as optimal transport, to provide a more detailed characterization of parameter similarities beyond the typical Euclidean distance.
-Additionally, further modularization of LoRA could enhance various applications, particularly in federated learning where strategies to minimize communication overhead and expedite model convergence through sharing and aggregating MSUs might be explored.
-We believe these advancements can significantly benefit a wide range of fields and applications.
+Potential future work includes exploring alternative distance metrics for LoRA-LEGO, such as optimal transport, to better characterize parameter similarities beyond the standard Euclidean distance.
+Additionally, further modularization of LoRA could enhance various applications.
+For example, in federated learning, strategies to minimize communication overhead and expedite model convergence through sharing and aggregating MSUs could be explored.
+We believe these advancements could significantly benefit a wide range of fields and applications.
 Appendix A Difference between LoAR Merging Setting and Model Merging Setting
@@
 In many cases, the goal of LoRA merging is to create a new LoRA that consolidates the capabilities of all involved LoRAs for simplified task-specific usage.
-In contrast, the LoRA merging approach presented in this paper bypasses this step; it directly merges the LoRA parameters to generate a unified LoRA with enhanced capabilities.
+In contrast, the LoRA merging setting presented in this paper bypasses the LoRA reload step; it directly merges the LoRA parameters to construct a unified LoRA with comprehensive capabilities.
 Appendix B Connection with Vanilla LoRA Composition Methods
-The vanilla LoRA composition comprises two training-free methods: the Ensemble of LoRAs and the Fusion of LoRAs.
-The Ensemble of LoRAs strategy involves the aggregation of the outputs of each submodule within the assembled LoRAs.
-Let us denote ${\mathbb{A}}=\{{\bm{A}}_{1},{\bm{A}}_{2},\ldots,{\bm{A}}_{n}\}$ and ${\mathbb{B}}=\{{\bm{B}}_{1},{\bm{B}}_{2},\ldots,{\bm{B}}_{n}\}$ as the sets representing submodules within $n$ LoRAs.
+The vanilla LoRA composition can be categoried into two types of training-free methods: the model ensembling and model merging Tang et al.
+(2024).
+The ensemble strategy involves aggregating the outputs of each submodule within the assembled LoRAs.
+Let us denote $\mathcal{A}=\{{\bm{A}}_{1},{\bm{A}}_{2},\ldots,{\bm{A}}_{n}\}$ and =$\mathcal{B}=\{{\bm{B}}_{1},{\bm{B}}_{2},\ldots,{\bm{B}}_{n}\}$ as the sets representing submodules within $n$ LoRAs.
 For an input ${\bm{x}}_{i}$, the output derived from the ensemble of LoRAs can be expressed as ${\bm{x}}_{i}^{\prime}=\frac{1}{n}\sum_{j=1}^{n}{\bm{B}}_{j}{\bm{A}}_{j}{\bm{x}}_{i}$, where ${\bm{x}}_{i}^{\prime}$ denotes the output.
-In contrast to the Ensemble method, which combines the output of different LoRAs, model merging presents an alternative composition strategy.
-Let the parameters of each LoRA, $\phi_{i}$, be denoted by $\Theta_{i}$.
-A typical strategy involves averaging these parameters, represented as $\Theta_{\text{fusion}}=\frac{1}{k}\sum_{j=1}^{k}\Theta_{j}$.
-This formulation allows the fused parameter to function akin to a single LoRA.
-Our proposed LoRA-LEGO method offers refined control over model complexity and performance, facilitating an optimal balance between semantic integration and computational efficiency.
-Unlike Simple Fusion, which lacks alignment and may introduce inconsistencies across merged LoRAs, our approach aligns and selectively fuses MSUs based on their semantic closeness.
-This selective fusion significantly reduces redundancy, a prevalent issue in the Ensemble method where all MSUs are treated uniformly, potentially leading to an excessive parameter count that can degrade performance.
-Our clustering-based strategy addresses these challenges by effectively condensing the most relevant semantic features into fewer clusters, thereby enhancing both the model’s performance and its inference efficiency.
+The performance of the ensemble of LoRAs tends to be more stable, but it incurs additional computational overhead.
+In contrast to the ensemble method, model merging presents an alternative composition strategy.
+A typical strategy involves employing element-wise fusion of these parameters, represented as ${\bm{A}}^{\prime}=\frac{1}{n}\sum_{j=1}^{n}{\bm{A}}_{j}$ and ${\bm{B}}^{\prime}=\frac{1}{n}\sum_{j=1}^{n}{\bm{B}}_{j}$.
+This formulation allows the merged parameters to function similarly to a single LoRA.
+However, directly merging parameters can lead to performance degradation due to parameter interference.
+Our proposed LoRA-LEGO method serves as a bridge between the two strategies, ensuring an optimal balance between computational efficiency and performance.
+By selectively aligning and fusing MSUs based on their semantic similarity, LoRA-LEGO effectively condenses the most relevant semantic features into fewer clusters.
+This process allows for the merging of parameters within each cluster, reducing the overall parameter count in a manner similar to the model merging method.
+By adjusting the number of clusters, LoRA-LEGO can accommodate more parameters for inference, much like the ensemble method.
+In this way, our method leverages the strengths of both methodologies, ultimately enhancing model performance and inference efficiency.
 Appendix C Details of Baseline Methods
@@
     Weight Averaging.
-This approach averages the parameters across different instances of LoRA, resulting in a new composite LoRA defined as ${\bm{A}}^{\prime}=\frac{1}{n}\sum_{i=1}^{n}{\bm{A}}_{i}$ and ${\bm{B}}^{\prime}=\frac{1}{n}\sum_{i=1}^{n}{\bm{B}}_{i}$, where ${\bm{A}}_{i}$ and $B_{i}$ represent the parameters from the $i$-th instance of the original LoRA models, and $n$ is the number of models being averaged.
+This approach averages the parameters across different instances of LoRA, resulting in a new composite LoRA defined as ${\bm{A}}^{\prime}=\frac{1}{n}\sum_{i=1}^{n}{\bm{A}}_{i}$ and ${\bm{B}}^{\prime}=\frac{1}{n}\sum_{i=1}^{n}{\bm{B}}_{i}$, where ${\bm{A}}_{i}$ and ${\bm{B}}_{i}$ represent the parameters from the $i$-th instance of the original LoRA models, and $n$ is the number of models being averaged.
   2. 2.